Find all the critical points of the function f(x) = x2 - 6x + 13. Explain how you arrive at your answer.
Answer #1

+2) = 2? 6x +13 diff writo f(x) = 2x-6 For Coutical Points f(x) = 0 2x-6:0 212-3) = 0 K-3-0 x= 3 Here witical of f(x) is x=